word problems involving quadratic equation
the area of a rectangle is 52 square centimeters. find its perimeter when the length is 1 cm more than 3 times its width.

Answer by nerdybill(7384)

the area of a rectangle is 52 square centimeters. find its perimeter when the length is 1 cm more than 3 times its width.
.
Let w = width
then
3w+1 = length
.
width*length = area
w(3w+1) = 52
3w^2+w = 52
3w^2+w-52 = 0
3w^2-12w+13w-52 = 0
(3w^2-12w)+(13w-52) = 0
3w(w-4) + 13(w-4) = 0
(w-4)(3w+13) = 0
w = {-13/3, 4}
toss out the negative value (extraneous) leaving:
w = 4 cm (width)
.
Length:
3w+1 = 3(4)+1 = 12+1 = 13 cm
.
Perimeter:
2(4+13) = 2(17) = 34 cm